OpenStudy (anonymous):

can somebody help with me with this 2/x+1+4/x-1/x(x+1) I need to solve it with x=2 I don't manage it

jhonyy9 (jhonyy9):

so there are 2 4 1 ----- + ----- - -------- = 0 so this is equal zero for being x+1 x x(x+1) one equation i think yes ?

jhonyy9 (jhonyy9):

and if there is x=2 like a root for this equation so than just you need substitute the 2 in place of x and make the calcule ok ?

jhonyy9 (jhonyy9):

wait or if you dont like cooperating nothing i will say bye

OpenStudy (anonymous):

I'm here

jhonyy9 (jhonyy9):

ok than make this calcule substitute the 2 in place of x and will get the result ok ?

OpenStudy (anonymous):

I wrote wrong it shall be 2/2+1+5/2-1/2(2+1)

OpenStudy (anonymous):

I've taken away all the x's now

OpenStudy (phi):

it would be good to use parentheses 2/2+1 is normally interpreted as \( \frac{2}{2}+1\) while 2/(2+1) means \( \frac{2}{2+1}\) But to do this problem, do the arithmetic. How far can you get?

OpenStudy (anonymous):

Add 1 to 2 to get 3. Multiply multiply To compute a product; to perform a multiplication. -1 by 3 to get -3. Reduce the expression expression A mathematical symbol, or combination of symbols, representing a value, or relation. Example: 2+2=4 by removing a factor factor One of two or more expressions that are multiplied together to get a product. of 2 from the numerator numerator The top part of a fraction. and denominator denominator The bottom part of a fraction. . Combine combine To join, or bring together. the numerators numerator The top part of a fraction. of all fractions fraction A number used to name a part of a group or a whole. The number below the division line is the denominator, and the number above the division line is the numerator. that have common denominators denominator The bottom part of a fraction. . Subtract 3 from 5 to get 2. Reduce the expression expression A mathematical symbol, or combination of symbols, representing a value, or relation. Example: 2+2=4 by removing a factor factor One of two or more expressions that are multiplied together to get a product. of 2 from the numerator numerator The top part of a fraction. and denominator denominator The bottom part of a fraction. . 1+1+1 Add 1 to 1 to get 2. 2+1 Add 1 to 2 to get 3. 3

OpenStudy (phi):

if the problem is \[ \frac{2}{x+1} +\frac{5}{x} -\frac{1}{x(x+1)}\]with x=2 \[ \frac{2}{2+1} +\frac{5}{2} -\frac{1}{2(2+1)}\] \[ \frac{2}{3} +\frac{5}{2} -\frac{1}{2\cdot (3)}\] \[ \frac{2}{3} +\frac{5}{2} -\frac{1}{6}\] you have 3 fractions, and need a common denominator. 6 will work multiply the first fraction by 2/2. multiply the 2nd fraction by 3/3 \[ \frac{2}{3}\cdot\frac{2}{2} +\frac{5}{2}\cdot\frac{3}{3} -\frac{1}{6}\] to multiply fractions, multiply top times top and bottom times bottom \[ \frac{4}{6} +\frac{15}{6} -\frac{1}{6}\] now add or subtract the numerators (tops) \[ \frac{4+15-1}{6} \] \[ \frac{18}{6} \] 3